OpenStudy (anonymous):

Please explain to me how to do B. I have done A. and got 10^-7 A. Medical X-rays, with a wavelength of about 10^-10m, can penetrate completely through your skin. Ultraviolet rays, which cause sun burn by penetrating only the top level of skin, have a wavelength about 1000 times the wavelength of an X-ray. find the wavelength of ultraviolet rays. Show your work. B. The wavelengths of visible light are between 3.8 * 10^-7m and 7.6 * 10^-7m. Are these wavelengths longer or shorter than those of ultraviolet rays? Explain

OpenStudy (goformit100):

is it a Mathematics Question ?

OpenStudy (anonymous):

Yes did you read it -_-

OpenStudy (anonymous):

I am doing my school right now I don't have time to read all of that sorry I will read it later

OpenStudy (goformit100):

Please Read CoC (compulsory to read for all "Open Study" users) : http://openstudy.com/code-of-conduct @connexus.com Sir, Please read it first before using Open Study

OpenStudy (anonymous):

I read it and I am not braking any rules :P

OpenStudy (goformit100):

it is a Physics Question Post it in Physics Section Please, otherwise other user will have difficulty.

OpenStudy (anonymous):

This is a math problem please re read it

OpenStudy (goformit100):

ok ... i thought it because it includes X-rays

OpenStudy (anonymous):

A lot of math problems involve science its ok

OpenStudy (kropot72):

If you put the wavelengths in fractional form the answer should be obvious: \[uv\ wavelength=\frac{1}{10^{7}}\] \[visible\ light\ wavelength\ range=\frac{3.8}{10^{7}} \to\ \frac{7.6}{10^{7}}\]

OpenStudy (goformit100):

I am really Sorry @connexus.com Sir. If possible please excuse me.

OpenStudy (anonymous):

@kropot72 I know that but I do not know what I am suppose to do for B

OpenStudy (kropot72):

Is 1/10^7 m longer or shorter than 3.8/10^7 m ?

OpenStudy (anonymous):

Shorter

OpenStudy (kropot72):

Correct! And is 1/10^7 m longer or shorter than 7.6/10^7 m ?

OpenStudy (anonymous):

Shorter

OpenStudy (kropot72):

Correct again. So you have found that the wavelength of ultra-violet rays is shorter than the wavelength of all colors in the spectrum of visible light.

OpenStudy (anonymous):

Ok so what would I put for B

OpenStudy (kropot72):

The wavelengths of visible light are longer than the wavelength of ultra-violet light.

OpenStudy (anonymous):

Thanks :)

OpenStudy (kropot72):

You're welcome :)